# D2USt3R: Enhancing 3D Reconstruction with 4D Pointmaps for Dynamic Scenes

### v1

@@ 4.2 Occlusion and dynamic masks preparation
To align dynamic objects from $I^{2}$ to $I^{1}$, we introduce a dynamic alignment loss that leverages both the occlusion mask ${M}_{\mathrm{occ}}$ and the dynamic mask ${M}_{\mathrm{dyn}}$ to effectively address occlusions and motion. As illustrated in Figure 4, these masks are computed in a dedicated pipeline. This method ensures that when points from the second view $I^{2}$, are transformed into the coordinate system of $I^{1}$, they accurately correspond to the temporal state of the first view. In line with our regression loss, we further incorporate confidence estimates to define a confidence-aware alignment loss. The dynamic alignment loss is formulated as follows:
$\mathcal{L}_{\text{dyn}}=$ (8)
$\sum_{i\in\mathcal{D}^{2}}M^{2}_{\mathrm{dyn},i}(1-M^{2}_{\mathrm{occ},i})C_{i}^{2,1}\left\|\frac{1}{\bar{z}_{1}}\bar{X}_{i+\mathbf{b}(i)}^{1,1}-\frac{1}{z_{1}}X_{i}^{2,1}\right\|$
$+\sum_{i\in\mathcal{D}^{1}}M^{1}_{\mathrm{dyn},i}(1-M^{1}_{\mathrm{occ},i})C_{i}^{1,2}\left\|\frac{1}{\bar{z}_{2}}\bar{X}_{i+\mathbf{f}(i)}^{2,2}-\frac{1}{z_{2}}X_{i}^{1,2}\right\|.$
We leverage optical flow $\mathbf{f}$ to establish dense correspondences between $I^{1}$ and $I^{2}$, and its reverse direction $\mathbf{b}$. This mechanism ensures that both views are aligned within a unified coordinate framework. The first term in our loss function enforces the alignment of the pointmap in the coordinate space of the camera system associated with $I^{1}$, while the second term introduces a symmetric constraint by aligning the points when the roles of the views are swapped.
Figure 5: Qualitative results for pointmap reconstruction. We qualitatively compare the 3D pointmap of D2USt3R against other pointmap regression models. All visualizations presents per-pixel pointmaps without applying confidence thresholding. It is notable that both DUSt3R [42] and MonST3R [48] struggle to accurately reconstruct scenes that include dynamic elements. We find that inaccurately established correspondence fields between dynamic regions negatively affect the overall reconstruction performance.
Figure 6: Depth estimation qualitative results.
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{static}}+\mathcal{L}_{\text{dyn}}$ (9)
@@ 4.4 Additional heads for downstream task
Since our model implicitly understands the mask’s location, we further regress a dynamic mask using an additional head. We predict a single-channel logit, $\hat{M}_{\text{dyn}}$, using a DPT head [31] in a manner similar to the pointmap regression head. To train this head, we employ the following cross-entropy loss:
$\mathcal{L}_{\text{mask}}=-\frac{1}{|\mathcal{D}_{\text{all}}|}$ $\sum_{i\in\mathcal{D}_{\text{all}}}\left[M_{\text{dyn},i}\log(\sigma(\hat{M}_{\text{dyn},i}))\right.$ (10)
$+\left.(1-M_{\text{dyn},i})\log(1-\sigma(\hat{M}_{\text{dyn},i}))\right],$
@@
For accurate optical flow estimation, we incorporate an additional flow head built on the RAFT architecture [39]. Inspired by insights from ZeroCo [1], our flow head utilizes cross-attention maps rather than conventional 4D correlation volumes. Empirically, we find that this helps to further enhances the accuracy of estiamted flow maps. We train this head using the Mixture-of-Laplace loss [44] as the supervision signal.
@@ 5 Experiments
Single-frame depth DepthAnythingv2 [46] 0.098 89.0 - - 0.073 93.8 - - 0.336 55.6 - - 0.069 93.7
Marigold [19] 0.205 72.3 - - 0.066 96.4 - - 0.623 50.5 - - 0.104 89.9
Multi-frame depth DUSt3R [42] 0.176 76.5 0.221 71.3 0.135 82.4 0.127 83.7 0.370 58.5 0.672 54.9 0.076 93.6
MASt3R [21] 0.165 79.0 0.199 73.8 0.183 77.5 0.167 79.5 0.330 57.3 0.528 54.4 0.050 96.8
MonST3R [48] 0.145 81.2 0.152 79.2 0.068 94.4 0.066 94.9 0.345 56.2 0.525 46.9 0.070 95.0
MonST3R* [48] 0.159 81.0 0.181 76.5 0.076 93.9 0.071 94.4 0.349 52.5 0.565 36.9 0.103 90.9
@@
DUSt3R [42] 0.141 82.5 0.424 58.7 0.112 86.3 0.080 90.7 0.176 76.8
MASt3R [21] 0.142 82.0 0.354 57.9 0.076 93.2 0.129 84.9 0.160 78.7
MonST3R [48] 0.076 93.9 0.345 56.5 0.101 89.3 0.091 88.8 0.147 81.1
MonST3R* [48] 0.083 93.6 0.387 50.6 0.143 85.0 0.084 90.1 0.163 79.1
@@
Table 4: Single-frame depth estimation results. *: Reproduced with same dataset as Ours.
@@
[…]
DUSt3R [42] 6.15 4.51 0.29 0.26 2.36 0.98 0.013 0.01 0.74 0.54 0.11 0.08
MASt3R [21] 4.71 3.40 0.23 0.19 2.83 1.13 0.06 0.03 0.85 0.64 0.05 0.04
MonST3R [48] 4.90 2.30 0.26 0.22 1.88 1.39 0.019 0.01 0.94 0.79 0.10 0.08
MonST3R* [48] 8.50 2.61 0.27 0.23 1.76 1.40 0.02 0.01 0.74 0.58 0.10 0.08
@@
Table 5: Camera pose estimation results. *: Reproduced with same dataset as Ours.
@@
DUSt3R [42] 30.96 35.11 14.19
MASt3R [21] 39.37 39.50 13.27
MonST3R [48] 38.47 41.92 14.91
MonST3R* [48] 37.47 40.58 14.58
@@
Croco-Flow [45] 3.31 4.28 13.24
SEA-RAFT [44] 5.21 13.18 4.43
@@
Table 6: Evaluation of pointmap alignment accuracy in dynamic objects. We evaluate the Flow End-Point Error (EPE) $\downarrow$ on the Sintel and KITTI datasets. Note that Croco-Flow [45] was grayed out in the Sintel evaluation since it was trained on the Sintel dataset. *: Reproduced with same dataset as Ours.
5.1 Experimental setup
Implementation details.
Building on top of DUSt3R [42], we freeze the encoder and fine-tune only the decoder and the DPT head [31], as done similarly by MonST3R. For each epoch, we randomly sample 20,000 image pairs and the network is trained for 50 epochs. We use the AdamW optimizer [26] with an initial learning rate of 5e-5. We train with 4 NVIDIA RTX 6000 GPUs, with a batch size of 4 images per GPU and gradient accumulation steps set to 2. More details can be found in .
@@ 5.1 Experimental setup
As shown in Table 2, we train D2USt3R  on multiple datasets, including BlinkVision Outdoor [22], BlinkVision Indoor [22], Spring [28], PointOdyssey [51], and TartanAir [43]. Each epoch consisted of sampling 7,750, 4,750, 2,500, 2,500, and 2,500 pairs, respectively. Additionally, we perform random sampling with temporal strides varying from 1 to 9 to account for large camera motions and dynamic scenarios.
@@
Following [42], we evaluate our method on depth estimation and camera pose estimation. We additionally evaluate pointmap alignment accuracy in dynamic regions. We compare our method against existing state-of-the-art pointmap regression models, specifically DUSt3R [42], MASt3R [21], and MonST3R [48]. Furthermore, to ensure fair comparisons and demonstrate the effectiveness of our approach for predicting the 4D pointmap, we trained a variant of MonST3R, termed MonST3R*, under the same setup as ours.
@@
For multi-frame depth estimation, we evaluate on the TUM-Dynamics [37], Bonn [29], Sintel [2], KITTI [11], and ScanNet [7] datasets using image pairs with source frames offset from the target by strides of 1, 3, 5, 7, and 9 frames. We assess performance over the entire scene and exclusively on dynamic regions when dynamic masks are available. For single-frame depth estimation, we evaluate on the Bonn [29], Sintel [2], KITTI [10], NYU-v2 [35], and TUM-Dynamics [37] datasets. In both settings, we follow the affine-invariant depth evaluation protocol, reporting the Absolute Relative Error (AbsRel) and the percentage of inlier points ($\delta_{1}$, where $\delta<1.25$).
@@ 5.2 Experimental results
In this experiment, we evaluate our method and compare with existing methods on single/multi-frame depth estimation. The results are summarized in Table 3 and  4. We also show qualitative comparisons in Figure 5 and Figure 6. From the results, we observe that D2USt3R outperforms other methods overall. However, our performance on KITTI is somewhat limited, likely due to the absence of driving scenes in our training data, which may have disadvantaged our model. Nonetheless, we compare with MonST3R, which was trained with the same training datasets as ours, and find that the performance is comparable. Finally, we highlight in Figure 5 that our approach consistently predicts accurate depth for dynamic human subjects.
@@
In Table 5, we evaluate camera pose performance of our model on the Sintel [2], TUM-dynamics [37] and ScanNet [7] datasets. Note that we use the 2D–3D matching between the prediction $X^{2,1}$ and the original coordinates to obtain camera poses via PNP-RANSAC.
@@
In Table 6, we evaluate pointmap alignment accuracy in dynamic objects. For this, we use Sintel [2] and KITTI [10]. Since we directly obtain aligned pointmaps even in dynamic objects, we can easily derive 2D-2D matching points by computing the difference between $X^{2,1}$ and $X^{1,1}$. From the results, we find that our method outperforms other baselines, which is further supported in Figure 7, where accurately captured correspondences between objects in different time step frames are observed. This gap is further broadened when we leverage an additional flow head, as shown in Table 6, where ours with the flow head outperforms the state-of-the-art SEA-RAFT [44]. We provide qualitative examples in Figure 8.
@@ 5.3 Ablation study
While we have already demonstrated the effectiveness of our proposed 3D alignment loss and 4D pointmaps are validated in previous experiments, we additionally explore the impacts of different finetuning strategies. The results are shown in Table 7. From the results, compared to full fine-tuning, partial fine-tuning of only the decoder and the downstream head provides more competitive results. We thus only train the decoder and head in this work.
Figure 8: Visualization of predicted optical flow.
Figure 9: Visualization of predicted dynamic mask.
@@ 6 Conclusion
Supplementary Material
@@ A. Training dataset
Table 8: List of excluded scenes from the datasets.
We excluded invalid scenes from the PointOdyssey, Blinkvision Outdoor utilized in Table 2. Specifically, in the PointOdyssey dataset, we removed scenes containing smoke or inaccurate annotations. Additionally, in the Blinkvision Outdoor dataset, we excluded scenes with incomplete dataset annotations. The excluded scenes are listed in Table 8.

### v2

@@ 4.3 Objective Function
To align dynamic region from $I^{2}$ to $I^{1}$, we introduce a dynamic alignment loss that leverages both the occlusion mask ${M}_{\mathrm{occ}}$ and the dynamic mask ${M}_{\mathrm{dyn}}$ to effectively address occlusions and motion. As illustrated in Figure 4, these masks are computed in a dedicated pipeline. This method ensures that when points from the second view $I^{2}$, are transformed into the coordinate system of $I^{1}$, they accurately correspond to the temporal state of the first view. In line with our regression loss, we further incorporate confidence estimates to define a confidence-aware alignment loss. The dynamic alignment loss is formulated as follows:
$\mathcal{L}_{\text{dyn}}=\sum_{i\in\mathcal{D}^{2}}M^{2}_{\mathrm{dyn},i}(1-M^{2}_{\mathrm{occ},i})C_{i}^{2,1}\left\|\frac{1}{\bar{z}_{1}}\bar{X}_{i+\mathbf{b}(i)}^{1,1}-\frac{1}{z_{1}}X_{i}^{2,1}\right\|-\alpha\log C_{i}^{2,1}$ (8)
$+\sum_{i\in\mathcal{D}^{1}}M^{1}_{\mathrm{dyn},i}(1-M^{1}_{\mathrm{occ},i})C_{i}^{1,2}\left\|\frac{1}{\bar{z}_{2}}\bar{X}_{i+\mathbf{f}(i)}^{2,2}-\frac{1}{z_{2}}X_{i}^{1,2}\right\|-\alpha\log C_{i}^{1,2}.$
We leverage optical flow $\mathbf{f}$ to establish dense correspondences between $I^{1}$ and $I^{2}$, and its reverse direction $\mathbf{b}$. The first term in our loss function enforces the alignment of the pointmap in the coordinate space of the camera system associated with $I^{1}$, while the second term introduces a symmetric constraint by aligning the points when the roles of the views are swapped.
Figure 5: Pointmap reconstruction. We qualitatively compare the 3D pointmap of D2USt3R against other pointmap regression models wang2024dust3r ; zhang2024monst3r . All visualizations presents per-pixel pointmaps without applying confidence thresholding. It is notable that both DUSt3R wang2024dust3r and MonST3R zhang2024monst3r struggle to accurately reconstruct scenes that include dynamic object. We find that inaccurately established correspondence fields between dynamic regions negatively affect the overall reconstruction performance.
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{static}}+\mathcal{L}_{\text{dyn}}.$ (9)
@@ 4.4 Additional heads for downstream task
To further enhance the capabilities and interpretability of our SDAP framework, we introduce two additional downstream heads dedicated to explicitly modeling dynamic masks and optical flow.
@@
Since our model implicitly encodes regions corresponding to dynamic motion, we explicitly regress dynamic masks using an additional head. Specifically, we predict a single-channel logit map, $\hat{M}_{\text{dyn}}$, using a DPT head ranftl2021vision , structured similarly to our pointmap regression head. We supervise this head using a binary cross-entropy loss defined as follows:
$\mathcal{L}_{\text{mask}}=-\frac{1}{|\mathcal{D}_{\text{all}}|}$ $\sum_{i\in\mathcal{D}_{\text{all}}}\left[M_{\text{dyn},i}\log(\sigma(\hat{M}_{\text{dyn},i}))+(1-M_{\text{dyn},i})\log(1-\sigma(\hat{M}_{\text{dyn},i}))\right],$ (10)
@@
To accurately estimate optical flow, we incorporate an additional head based on the RAFT architecture teed2020raft . Inspired by recent findings from ZeroCo an2024cross , our flow head utilizes cross-attention maps instead of traditional 4D correlation volumes. The optical flow head is supervised using the Mixture-of-Laplace loss wang2024sea .
@@ 5 Experiments
5.1 Experimental setup
Implementation details.
Building on top of DUSt3R wang2024dust3r , we freeze the encoder and fine-tune only the decoder and the DPT head ranftl2021vision , as done similarly by MonST3R. For each epoch, we randomly sample 20,000 image pairs and the network is trained for 50 epochs. We use the AdamW optimizer loshchilov2017decoupled with an initial learning rate of 5e-5. We train with 4 NVIDIA RTX 6000 GPUs, with a batch size of 4 images per GPU and gradient accumulation steps set to 2.
@@ 5.1 Experimental setup
Single-frame depth DepthAnythingv2 depth_anything_v2 0.098 89.0 - - 0.073 93.8 - - 0.336 55.6 - - 0.069 93.7
Marigold ke2023repurposing 0.205 72.3 - - 0.066 96.4 - - 0.623 50.5 - - 0.104 89.9
Multi-frame depth DUSt3R wang2024dust3r 0.176 76.5 0.221 71.3 0.135 82.4 0.127 83.7 0.370 58.5 0.672 54.9 0.076 93.6
MASt3R leroy2024grounding 0.165 79.0 0.199 73.8 0.183 77.5 0.167 79.5 0.330 57.3 0.528 54.4 0.050 96.8
MonST3R zhang2024monst3r 0.145 81.2 0.152 79.2 0.068 94.4 0.066 94.9 0.345 56.2 0.525 46.9 0.070 95.0
MonST3R* zhang2024monst3r 0.159 81.0 0.181 76.5 0.076 93.9 0.071 94.4 0.349 52.5 0.565 36.9 0.103 90.9
@@
Table 3: Single-frame depth estimation results. *: Reproduced using the same dataset as ours.
@@
DUSt3R wang2024dust3r 0.141 82.5 0.424 58.7 0.112 86.3 0.080 90.7 0.176 76.8
MASt3R leroy2024grounding 0.142 82.0 0.354 57.9 0.076 93.2 0.129 84.9 0.160 78.7
MonST3R zhang2024monst3r 0.076 93.9 0.345 56.5 0.101 89.3 0.091 88.8 0.147 81.1
MonST3R* zhang2024monst3r 0.083 93.6 0.387 50.6 0.143 85.0 0.084 90.1 0.163 79.1
@@
Table 4: Camera pose estimation results. *: Reproduced using the same dataset as ours.
@@
$downarrow Avg\downarrow Med\downarrow$ $downarrow Avg\downarrow Med\downarrow$ $downarrow Avg\downarrow Med\downarrow$
DUSt3R wang2024dust3r 6.15 4.51 0.29 0.26 2.36 0.98 0.013 0.01 0.74 0.54 0.11 0.08
MASt3R leroy2024grounding 4.71 3.40 0.23 0.19 2.83 1.13 0.06 0.03 0.85 0.64 0.05 0.04
MonST3R zhang2024monst3r 4.90 2.30 0.26 0.22 1.88 1.39 0.019 0.01 0.94 0.79 0.10 0.08
MonST3R* zhang2024monst3r 8.50 2.61 0.27 0.23 1.76 1.40 0.02 0.01 0.74 0.58 0.10 0.08
@@
Table 5: Pointmap alignment accuracy in dynamic objects. We report End-Point Error (EPE) $\downarrow$ on the Sintel and KITTI datasets. Note that Croco-Flow weinzaepfel2023croco is shown in gray on the Sintel benchmark because it was trained on that dataset. *: Reproduced using the same dataset as ours.
@@
DUSt3R wang2024dust3r 30.96 35.11 14.19
MASt3R leroy2024grounding 39.37 39.50 13.27
MonST3R zhang2024monst3r 38.47 41.92 14.91
MonST3R* zhang2024monst3r 37.47 40.58 14.58
@@
Croco-Flow weinzaepfel2023croco 3.31 4.28 13.24
SEA-RAFT wang2024sea 5.21 13.18 4.43
@@
Figure 6: Depth estimation qualitative results.
@@
As shown in Table 1, we train D2USt3R on multiple datasets, including BlinkVision Outdoor li2024blinkvision , BlinkVision Indoor li2024blinkvision , Spring mehl2023spring , PointOdyssey zheng2023pointodyssey , and TartanAir wang2020tartanair . Each epoch consisted of sampling 7,750, 4,750, 2,500, 2,500, and 2,500 pairs, respectively. Additionally, we perform random sampling with temporal strides varying from 1 to 9 to account for large camera motions and dynamic scenarios.
@@
Following wang2024dust3r , we evaluate our method on depth and camera pose estimation. We additionally evaluate pointmap alignment accuracy in dynamic regions. We compare our method against existing state-of-the-art pointmap regression models, specifically DUSt3R wang2024dust3r , MASt3R leroy2024grounding , and MonST3R zhang2024monst3r . Furthermore, to ensure fair comparisons and demonstrate the effectiveness of our approach, we trained a variant of MonST3R, termed MonST3R*, under the same setup as ours.
@@
For multi-frame depth estimation, we evaluate on the TUM-Dynamics sturm2012benchmark , Bonn palazzolo2019refusion , Sintel butler2012naturalistic , KITTI geiger2013vision , and ScanNet dai2017scannet datasets using image pairs with source frames offset from the target by strides of 1, 3, 5, 7, and 9 frames. We assess performance over the entire scene and exclusively on dynamic regions when dynamic masks are available. For single-frame depth estimation, we evaluate on the Bonn palazzolo2019refusion , Sintel butler2012naturalistic , KITTI geiger20133d , NYU-v2 silberman2012indoor , and TUM-Dynamics sturm2012benchmark datasets. In both settings, we follow the affine-invariant depth evaluation protocol, reporting the Absolute Relative Error (AbsRel) and the percentage of inlier points ($\delta_{1}$, where $\delta<1.25$).
Figure 8: Predicted optical flow.
Figure 9: Predicted dynamic mask.
@@ 5.2 Experimental results
In this experiment, we evaluate our method and compare with existing methods on single/multi-frame depth estimation. The results are summarized in Table 2 and  3. We also show qualitative comparisons in Figure 5 and Figure 6. From the results, we observe that D2USt3R outperforms other methods overall. However, our performance on KITTI is somewhat limited, likely due to the absence of driving scenes in our training data, which may have disadvantaged our model. Nonetheless, we compare with MonST3R, which was trained with the same training datasets as ours, and find that the performance is comparable. Finally, we highlight in Figure 5 that our approach consistently predicts accurate depth for dynamic human subjects.
@@
In Table 4, we evaluate camera pose performance of our model on the Sintel butler2012naturalistic , TUM-dynamics sturm2012benchmark and ScanNet dai2017scannet datasets. Note that we use the 2D–3D matching between the prediction $X^{2,1}$ and the original coordinates to obtain camera poses via PNP-RANSAC.
@@
In Table 5, we evaluate pointmap alignment accuracy in dynamic objects. For this, we use Sintel butler2012naturalistic and KITTI geiger20133d . Since we directly obtain aligned pointmaps even in dynamic objects, we can easily derive 2D-2D matching points by computing the difference between $X^{2,1}$ and $X^{1,1}$. From the results, we find that our method outperforms other baselines, which is further supported in Figure 7, where accurately captured correspondences between objects in different time step frames are observed. This gap is further broadened when we leverage an additional flow head, as shown in Table 5, where ours with the flow head outperforms the state-of-the-art SEA-RAFT wang2024sea . We provide qualitative examples in Figure 8.
@@
5.3 Robustness analysis on frame intervals
Table 6: Robustness analysis on frame intervals. We evaluate the robustness of our method with respect to varying frame intervals $\Delta t\in\{1,3,5,7,9\}$. *: Reproduced using the same dataset as ours.
Methods $\Delta t=1$ $\Delta t=3$ $\Delta t=5$ $\Delta t=7$ $\Delta t=9$
[…]
MonST3R*zhang2024monst3r 0.078 93.8 0.078 93.5 0.075 93.9 0.074 93.9 0.072 94.3
D2USt3R(Ours) 0.061 95.6 0.061 95.5 0.060 95.7 0.061 95.9 0.058 96.2
We evaluate multi-frame depth estimation for the robustness of our method with respect to varying frame intervals on the Bonn dataset palazzolo2019refusion , as summarized in Table 6. Across all tested intervals ($\Delta t\in\{1,3,5,7,9\}$), the results show that our model maintains stable performance and even exhibits slight improvement as the interval widens, demonstrating strong temporal consistency and resilience to larger motion baselines. These results suggest that the proposed flow-chaining strategy effectively preserves geometric coherence across diverse motion dynamics, enabling reliable depth estimation under varying temporal conditions.
Table 7: Ablation on training strategy.
@@ 5.4 Ablation study
Although we’ve already validated our 3D alignment loss and effectiveness of proposed SDAP, we further examine how different fine-tuning strategies affect performance. Since our method, unlike DUSt3R or MonST3R, learns precise maps of cross-attention (Figure 3), we ask whether updating the encoder along with the decoder and head yields gains. As shown in Table 7, fine-tuning only the decoder and downstream head actually outperforms full fine-tuning. Therefore, in this work we restrict training to the decoder and head.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This research was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, RS-2024-00509279, RS-2025-II212068, RS-2023-00227592, RS-2025-02214479, RS-2024-00457882, RS-2025-25441838, RS-2025-25441838, RS-2025-02214479, RS-2025-02217259) and the Culture, Sports, and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism (RS-2024-00345025, RS-2024-00333068, RS-2023-00222280, RS-2023-00266509), and National Research Foundation of Korea (RS-2024-00346597).
This research was supported by the ETH AI Center through an ETH AI Center postdoctoral fellowship to Sunghwan Hong.
@@ A. Training dataset
Table 8: List of excluded scenes from the datasets.
@@
We provide the scene names that were excluded during our training. These scenes are excluded, either because their annotations include errors. Please refer to Table 8.